OpenStudy (stormoshawty):

Solve 7x-3y=20 5x+3y=16 using substitution.

OpenStudy (shamil98):

You don't need substitution for this. Just add the equations together. 7x - 3y = 20 5x + 3y = 16 the y's cross out. you get 12x = 36 x = 3 sub x into original equation to get the value of y.

OpenStudy (shamil98):

if you had to use substitution you could 7x - 20 = 3y 7/3x - 20/3 = y but it's fractions and using elimination you get whole numbers ( a whole lot neater, and less tedious)

OpenStudy (stormoshawty):

I know, but I used elimination on it before, now I wanted to use substitution.

OpenStudy (shamil98):

Okay. then sub for y with the given 7/3 x - 20/3 = y 5x + 3(7/3x - 20/3) = 16 just a hint since everything is multiplied by 3 and there are three in the denominators they cancel leaving just the top left. so 5x + 7x - 20 = 16

OpenStudy (stormoshawty):

x = 3

OpenStudy (shamil98):

yep.

OpenStudy (stormoshawty):

How do i solve for y now? Walk me through this.

OpenStudy (stormoshawty):

Nevermind. I'll just do the adding. How do i solve for y using addition?

OpenStudy (shamil98):

addition?

OpenStudy (shamil98):

you mean elimination?

OpenStudy (shamil98):

you're making things complicated if you wanted to solve for y first. 7x-3y=20 5x+3y=16 it would be best to solve for x and sub the value of x in and then find y.

OpenStudy (stormoshawty):

Lol. Okay, fine. How do I solve this with substitution.

OpenStudy (stormoshawty):

Shammmy.

OpenStudy (shamil98):

solve what?

OpenStudy (stormoshawty):

I want to get Y.

OpenStudy (shamil98):

you already your x value which is 3. so 7(3) - 3y = 20 evaluate and solve for y

OpenStudy (stormoshawty):

\[y = \frac{ 1 }{ 3 }\]

OpenStudy (shamil98):

yep